paul is seven years younger than his brother. in four years, his brother will be twice as old as paul. how old is his brother now?
Answer by majo89(3)

paul's age,p=brother's age(b)-7
in 4 years
b+4=2(p+4)
b-2p=4;
b-2(b-7)=4
b-2b=4-14
b=10 years